# Superquantile-Gibbs Relaxation for Minima-selection in Bilevel Optimization

### v2

@@ Abstract
Bilevel optimization (BLO) becomes fundamentally more challenging when the lower-level objective admits multiple minimizers. In contrast to the commonly studied unique-minimizer setting, this regime introduces two additional difficulties: (1) evaluating the hyper-objective $F_{\max}$ requires a minima-selection step, which, in the worst case, amounts to solving an optimization problem over a topologically disconnected set; (2) $F_{\max}$ is in general discontinuous without strong structural assumptions.
In this paper, we show that both difficulties can be circumvented under a local Polyak–Łojasiewicz (PŁ) property, denoted by , on the lower-level objective. This condition is strictly weaker than the global PŁ property assumed in recent BLO studies, encompassing a broader range of problems, such as hyperparameter tuning in over-parameterized deep learning. The implications of the condition are two-folded: Under this condition, we first show that $F_{\max}$ is Lipschitz continuous; Moreover, we identify that, for all upper-level variables, the corresponding set of lower-level minimizers are topologically connected and are provably closed embedded submanifolds, all sharing a common intrinsic dimension k.
Importantly, we show that this intrinsic dimension k, rather than the ambient dimension, fundamentally governs the complexity of BLO: We propose a method that finds an $(\epsilon,\rho)$-Goldstein stationary point of $F_{\max}$ using at most $\mathcal{O}\big(m^{8\texttt{k}+9}(\epsilon\rho)^{-8\texttt{k}-10}\big)$ queries to the gradient oracle of the lower-level function, where $m$ is the dimension of the upper-level domain. The key ingredient behind this result is a novel construction of an approximate function-evaluation oracle for $F_{\max}$: via a Superquantile–Gibbs relaxation, we reduce the minima-selection step, an optimization problem over a non-convex set, to a sampling problem that can be efficiently addressed using standard Langevin dynamics. To the best of our knowledge, this is the first work that rigorously tackles minima selection in BLO and explicitly characterizes how its complexity scales with the intrinsic dimensionality of the lower-level problem.
Keywords: Bilevel optimization, non-convex lower-level problem, and local PŁ condition.
@@ 1 Introduction
We study the pessimistic/optimistic formulation of Bilevel Optimization (BLO), which is a common paradigm in the literature (ye1997exact; ye1997necessary; dempe2007new; liu2021towards; xu2014smoothing), although alternatives to pessimistic/optimistic formulation for BLO also exist (see e.g., (arbel2022non; jiang2025correspondence; bolte2025bilevel)). The pessimistic formulation of BLO is defined as follows [^1] All results extend to the optimistic setting where minimizing is considered instead of maximizing $f$ over $x$.:
@@
where $\theta$ and $x$ are the upper- and lower-level variables, $f$ and $g$ are the upper- and lower-level objectives, respectively, and $\Theta$ is a closed compact convex subset of $\mathbb{R}^{m}$. We focus on the setting where the lower-level minima set $\mathcal{S}(\theta)$ is non-singleton. We refer to the maximization of $f$ over $x\in\mathcal{S}(\theta)$ as minima-selection and call $F_{\max}$ the hyper-objective. BLO captures problems with a nested structure, such as hyper-parameter optimization (feurer2019hyperparameter), meta-learning (bertinetto2018meta), and reinforcement learning (hong2023two; liu2021investigating). In particular, when the lower-level problem pertains to deep learning, the set $\mathcal{S}(\theta)$ is typically nonconvex and non-singleton (lin2024exploring; kuditipudi2019explaining; nguyen2019connected; garipov2018loss; draxler2018essentially; liang2018understanding; venturi2018spurious; sagun2017empirical). Therefore, minimia-selection becomes important.
@@
Recent work on BLO have studied nonconvex lower-level problems under the global Polyak-Łojasiewicz (PŁ) condition (chen2023bilevel; kwon2023penalty; xiao2023generalized; huang2024optimal; shen2025penalty; liu2022bome). This condition relaxes the classical strong convexity assumption on $g$, which implies singleton $\mathcal{S}(\theta)$, to cover nonconvex and potentially non-singleton optimal solution sets. Another motivation for adopting the global PŁ condition is to guarantee the continuity of the hyper-objective function $F_{\max}$, which might otherwise become discontinuous and render BLO problems generally intractable (kwon2023penalty; chen2023bilevel).
All of the aforementioned work and their results are crucially dependent on the assumption that $g$ is $\mathcal{C}^{2}$ and satisfies the global PŁ condition. Under these conditions, it can be proved that the lower-level solution set $\mathcal{S}(\theta)$ would be limited to either a singleton or an unbounded subset of $\mathbb{R}^{d}$ (criscitiello2025).
More specifically, huang2024optimal assume a nondegenerate Hessian of $g$ at global minimizers, implicitly enforcing uniqueness; liu2022bome explicitly assume $\mathcal{S}(\theta)$ is a singleton; xiao2023generalized and shen2025penalty assume additional convexity of the lower-level objective, implying that the minimizer set is a linear subspace; and kwon2023penalty require $g$ to be coercive which again yields a singleton solution222kwon2023penalty assume the proximal error bound for $g$ (with $\sigma=0$), equivalent to the global PŁ condition for smooth $g$ (chen2023bilevel, Prop. C.1).. The above observations hint the necessity of a further relaxation to capture richer landscapes.
@@
We relax the global PŁ assumption to the condition, which was recently proposed by gong2024poincare, to accommodate applications of interest such as those in deep learning. For a fixed $\theta$, a lower-level objective $g(\theta,\cdot)$ admits a provably connected optimal set $\mathcal{S}(\theta)$ and it satisfies the PŁ inequality locally in a neighborhood of $\mathcal{S}(\theta)$. Moreover, the condition ensures that $\mathcal{S}(\theta)$ is a $\mathcal{C}^{2}$ embedding submanifold of the ambient space $\operatorname{\mathbb{R}}^{d}$ without boundary (gong2024poincare). We shall define condition formally in Section˜2. This setting is sufficiently rich as the lower-level problem admits bounded non-singleton minimizers. Hence, the minima-selection step remains relevant. One such example is when $g$ represents the loss function of some deep learning tasks in the over-parameterized regime, where the local minima are observed to be connected (draxler2018essentially; garipov2018loss; nguyen2019connected; kuditipudi2019explaining; lin2024exploring). Another advantage of this setting is that, as we prove in Section˜3.1, the hyper-objective under the condition is Lipschitz continuous. However, under the same condition, the hyper-objective function is not necessarily differentiable, and conditions on functions $f$ and $g$ that would ensure the differentiability of the hyper-objective are often restrictive and unrealistic in practice (see the discussion in Section˜1.4).
@@
Solving BLO problems with a non-singleton lower-level solution set $\mathcal{S}(\theta)$ involves two main challenges: 1) As previously emphasized, for a fixed upper-level variable $\theta$, the minima selection problem requires maximization over a nonconvex set. Consequently, exact evaluations of the hyper-objective $F_{\max}$ are generally inaccessible. 2) Even when the minima selection subproblem can be solved exactly, the resulting hyper-objective $F_{\max}$ may still be non-differentiable with respect to the upper-level variable $\theta$. Addressing this issue requires establishing local optimality guarantees for a nonsmooth nonconvex objective, a widely recognized challenge in the literature (kornowski2024hardness; jordan2023deterministic; tian2022finite; kornowski2021oracle; zhang2020complexity).
To address these issues, we introduce the superquantile–Gibbs relaxation to approximate the hyper-objective, providing a practical surrogate for $F_{\max}$. Using evaluations of this surrogate, we propose a zeroth-order method that seeks a generalized Goldstein stationary point of $F_{\max}$ (see Section˜5.2; cf. (liu2024zeroth)).
@@ 1.1 Finding a Goldstein Stationary Point of the Hyper-objective
As we shall see in ˜4.5, under the condition the hyper-objective $F_{\max}$ is Lipschitz continuous. A natural choice for approximate stationarity in this setting is Clarke’s notion of stationarity (clarke1990optimization). However, zhang2020complexity showed that, for nonsmooth Lipschitz problems, obtaining an $\epsilon$-Clark stationary point in polynomial time is generally impossible, motivating more nuanced concepts. We therefore adopt the notion of $(\epsilon,\rho)$-Goldstein stationarity (goldstein1977optimization).
For the domain-constrained objective in (1), akin to (liu2024zeroth), we measure stationarity by a generalized $(\epsilon,\rho)$-Goldstein criterion which is defined for the gradient mapping [^3] The gradient mapping for $F:\Theta\to\mathbb{R}$ at a point $\theta$ is $\mathcal{G}_{\Theta}(\theta,\nabla F;\eta):=\eta^{-1}(\theta-\text{Proj}_{\Theta}(\theta-\eta\nabla F(\theta)))$., denoted by $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point (see Section˜5.2). As $F_{\max}$ may not be differentiable, we propose a zeroth-order (ZO) method to find such a stationary point of $F_{\max}$.
@@ 1.2 Superquantile-Gibbs Relaxation (SQ-G)
In this part, we introduce the superquantile of the random variable $f(\theta,X)$ where $X$ is sampled according to the the Gibbs measure $\sim\exp(-g(\theta,x)/\lambda)$. Our idea is that as $\lambda\rightarrow 0$, the Gibbs measure concentrates on the feasible domain $\mathcal{S}(\theta)$, and the superquantile at a high confidence level would approach the maximum value of $f(\theta,\cdot)$ on $\mathcal{S}(\theta)$. Formally, we define the Superquantile-Gibbs Relaxation (SQ-G)-loss using the variational formulation of the superquantile by (rockafellar2000optimization):
@@ 1.3 Our contributions
    To allow for lower-level optimal sets with rich structures, we relax the global PŁ condition, commonly imposed on the lower-level objective, to the weaker condition from (gong2024poincare). We show that the condition is sufficient to ensure Lipschitz continuity of the hyper-objective function $F_{\max}$.
@@
    We use the Superquantile–Gibbs relaxation (SQ-G) to approximate $F_{\max}$. This converts the crucial minima-selection step in BLO—originally a nonconvex optimization over a nonconvex set—into sampling from a Gibbs distribution, which we implement efficiently via Langevin dynamics.
@@
    Assume that $g$ satisfies and the regularity conditions for $f$ and $g$ (to be stated in Section˜2). As shown in Section˜5, by using a Projected Stochastic Zeroth-Order method with minima selection (see Algorithm˜2) to optimize $F_{\max}$, we obtain an $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point of $F_{\max}$ on average using $\mathcal{O}\big(m^{2\texttt{k}+3}(\epsilon\rho)^{-2\texttt{k}-4}\big)$ samples from the distribution $\mu^{\lambda}_{\theta}$ for $\texttt{k}\geq 1$. When the samples of $\mu^{\lambda}_{\theta}$ are approximated generated by the Langevin Monte Carlo method, we show in Section˜5.3 that the same goal can be achieved with $\mathcal{O}\big(m^{8\texttt{k}+9}(\epsilon\rho)^{-8\texttt{k}-10}\big)$ queries to a gradient oracle for $g$.
@@
    We evaluate our approach on both synthetic and real-world benchmarks and demonstrate the effect of minima selection. As shown in Section˜6, our proposed approach outperforms hyper-gradient and penalty-based baselines.
@@
(Exponential dependence on k) The $\epsilon^{-\texttt{k}}$ factor in our results reflects the intrinsic difficulty of the problem. Since evaluating $F_{\max}(\theta)=\max_{y\in S(\theta)}f(\theta,y)$ is equivalent to solving a nonconvex optimization problem over a nonconvex set, a single $\epsilon$-accurate evaluation of $F_{\max}(\theta)$ yields an $\Omega(\epsilon^{-\texttt{k}})$ complexity lower bound. More specifically, consider the instance where $S(\theta)=\mathbb{S}^{\texttt{k}}$ is a hypersphere. Reparameterizing $y$ in polar coordinates yields $F_{\max}(\theta)=\max_{x\in\mathbb{B}_{\texttt{k}}}f(\theta,y(x))$. Here $\mathbb{B}_{\texttt{k}}=[0,\pi]^{\texttt{k}}$ and the point $y(x)\in\mathbb{S}^{\texttt{k}}$ is given by $y_{i}(x)=\sin(x_{1})\cdots\sin(x_{i-1})\cos(2x_{i})$ for $i\in[\texttt{k}]$ and $y_{\texttt{k}+1}(x)=\sin(x_{1})\cdots\sin(x_{\texttt{k}-1})\sin(2x_{\texttt{k}})$. It follows from (nemirovskij1983problem, Section 1.6.2) that the worst-case complexity of finding an $\epsilon$-accurate solution to is $\Omega(\epsilon^{-\texttt{k}})$ for $f(\theta,\cdot)\in\mathcal{C}^{1}$.
@@ 1.4 Related work
Bilevel optimization has a long history in mathematical programming; see, e.g., the early bibliographic review by vicente1994bilevel, the monograph dempe2002foundations, the survey colson2007overview, the classical text on mathematical programs with equilibrium constraints (MPECs) by luo1996mathematical, and the recent collection dempe2020bilevel. A large part of this literature treated bilevel problems by transforming them into single-level ones, either through value-function reformulations or by replacing the lower-level problem with its Karush–Kuhn–Tucker (KKT) optimality system. In the latter case, the KKT complementarity relations between multipliers and constraint violations appear explicitly as equilibrium constraints, so the resulting problem is a mathematical program with equilibrium constraints; see, for example, ye1995optimality; ye2010new; dempe2012kkt; zemkoho2021comparison. In many practical models, these KKT-based formulations were implemented using so-called big-$M$ techniques, where large constants were introduced to linearize complementarity conditions. Choosing suitable big-$M$ constants is delicate: if they are too small, feasible bilevel solutions may be lost, whereas overly large values can lead to weak relaxations and numerical instability. These issues were analyzed in detail by kleinert2020bigM.
@@
Modern large-scale machine-learning applications led to gradient-based methods for bilevel empirical risk minimization and hyperparameter optimization. When the lower-level objective is strongly convex in the inner variable and $f,g$ are smooth, the hyper-objective is differentiable and its gradient can be obtained by implicit differentiation. Representative works in this setting included approximate hyper-gradient methods such as HOAG (pedregosa2016hoag), stochastic bilevel schemes (ghadimi2018approx; yang2021provably), and complexity-refined algorithms for empirical risk minimization (dagreou2024lowerbound; chen2024optimal), as well as more abstract treatments of optimization layers via automatic differentiation (ablin2020super), amortized implicit differentiation (arbel2021amortized), and modular implicit differentiation frameworks (blondel2022efficient; grazzi2020iteration). In these approaches, computing exact hyper-gradients typically required Hessian or Hessian-inverse information for the lower-level problem, which could be prohibitive in large neural networks. This motivated a substantial line of Hessian-free or fully first-order methods that approximated the hyper-gradient using only first-order information, including BOME (liu2022bome), Hessian/Jacobian-free stochastic methods with improved sample complexity (yang2023hessianfree), fully first-order stochastic approximation methods such as F2BO (kwon2023fully), and bilevel stochastic-gradient methods that also accommodate nonlinear lower-level constraints and inexact hyper-gradient computations (giovannelli2023inexact). These developments made bilevel optimization scalable in settings with a unique lower-level solution, so the minima-selection problem that we study did not arise.
@@
A complementary line of work relaxed lower-level strong convexity to a global Polyak–Łojasiewicz (PŁ) condition and developed bilevel gradient methods under this assumption (chen2023bilevel; kwon2023penalty; xiao2023generalized; huang2024optimal; shen2025penalty; liu2022bome). To obtain a smooth and well-behaved hyper-objective, these methods typically combined the global PŁ condition with additional regularity assumptions on the lower-level problem (as discussed earlier), so that minima selection was effectively avoided. In contrast, our analysis is based on a local PŁ condition and explicitly exploits the geometry of nontrivial minimizer manifolds. In a complementary direction, chen2025setsmoothness assumes a set-smoothness property of the solution mapping (implied by uniform error-bound/PŁ-type conditions) and shows that the induced optimistic/pessimistic hyper-objectives are weakly convex/concave, enabling computation of approximate Clarke hyper-stationary points. In contrast, we work under a local condition that allows nontrivial minimizer manifolds and does not rely on weak convexity/concavity of the hyper-objective; instead, we handle non-uniqueness via a superquantile–Gibbs surrogate and derive dimension-aware complexity bounds driven by $\dim(\mathcal{S}(\theta))$.
Morse- and correspondence-based formulations.

### v4

@@ Abstract
Bilevel optimization (BLO) becomes fundamentally more challenging when the lower-level objective admits multiple minimizers. In contrast to the commonly studied unique-minimizer setting, this regime introduces two additional difficulties: (1) evaluating the hyper-objective $F_{\max}$ requires a minima-selection step, which, in the worst case, amounts to solving an optimization problem over a topologically disconnected set; (2) $F_{\max}$ is in general discontinuous without strong structural assumptions. We show that both issues can be circumvented under a local Polyak–Łojasiewicz (PŁ) property, denoted by , on the lower-level objective. This condition is strictly weaker than the global PŁ property assumed in recent BLO studies and covers problems such as hyperparameter tuning in over-parameterized deep learning. Under , we prove that $F_{\max}$ is Lipschitz continuous and that, for all upper-level variables, the corresponding lower-level minimizer sets are topologically connected and are provably closed embedded submanifolds, all sharing a common intrinsic dimension k. We show that this intrinsic dimension k, rather than the ambient dimension, governs the complexity of BLO: We propose a method that finds an $(\epsilon,\rho)$-Goldstein stationary point of $F_{\max}$ using at most $\tilde{\mathcal{O}}\big(m^{8\texttt{k}+11}\epsilon^{-2}(\epsilon\rho)^{-8\texttt{k}-10}\big)$ queries to the gradient oracle of the lower-level function, where $m$ is the dimension of the upper-level domain. The key ingredient is a Superquantile–Gibbs relaxation that yields an approximate function-evaluation oracle for $F_{\max}$: it turns minima selection into a sampling problem that can be efficiently addressed using standard Langevin dynamics. To the best of our knowledge, this is the first work that rigorously tackles the minima selection step in BLO and explicitly characterizes how the overall complexity of BLO scales with the intrinsic dimensionality of the lower-level problem.
Keywords: Bilevel optimization, Nonconvex lower-level problem, Local PŁ condition, and Nonsmooth optimization.
@@ 1 Introduction
We study the pessimistic/optimistic formulation of Bilevel Optimization (BLO), which is a common paradigm in the literature [90, 87, 23, 62, 84], although alternatives to pessimistic/optimistic formulation for BLO also exist (see e.g., [4, 43, 11]). The pessimistic formulation of BLO is defined as follows [^1] All results extend to the optimistic setting where minimizing is considered instead of maximizing $f$ over $x$.:
@@
where $\theta$ and $x$ are the upper- and lower-level variables, $f$ and $g$ are the upper- and lower-level objectives, respectively, and $\Theta$ is a closed compact convex subset of $\mathbb{R}^{m}$. We focus on the setting where the lower-level minima set $\mathcal{S}(\theta)$ is non-singleton. We refer to the maximization of $f$ over $x\in\mathcal{S}(\theta)$ as minima-selection and call $F_{\max}$ the hyper-objective. BLO captures problems with a nested structure, such as hyper-parameter optimization [28], meta-learning [8], and reinforcement learning [39, 61]. In particular, when the lower-level problem pertains to deep learning, the set $\mathcal{S}(\theta)$ is typically nonconvex and non-singleton [58, 48, 69, 31, 27, 55, 80, 75]. Therefore, minima-selection becomes important.
@@
Recent works on BLO have studied nonconvex lower-level problems under the global Polyak-Łojasiewicz (PŁ) condition [14, 50, 83, 40, 77, 59]. This condition relaxes the classical strong convexity assumption on $g$, which implies singleton $\mathcal{S}(\theta)$, to cover nonconvex and potentially non-singleton optimal solution sets. Another motivation for adopting the global PŁ condition is to guarantee the continuity of the hyper-objective function $F_{\max}$, which might otherwise become discontinuous and render BLO problems generally intractable [50, 14].
All of the aforementioned work and their results are crucially dependent on the assumption that $g$ is $\mathcal{C}^{2}$ and satisfies the global PŁ condition. Under these conditions, it can be proved that the lower-level solution set $\mathcal{S}(\theta)$ would be limited to either a singleton or an unbounded subset of $\mathbb{R}^{d}$ [20]. More specifically, Huang [40] assume a nondegenerate Hessian of $g$ at global minimizers, implicitly enforcing uniqueness; Liu et al. [59] explicitly assume $\mathcal{S}(\theta)$ is a singleton; Xiao et al. [83] and Shen et al. [77, Sec. 4] assume additional convexity of the lower-level objective, implying that the minimizer set is a linear subspace; and Kwon et al. [50] require $g$ to be coercive which again yields a singleton solution [^2] Kwon et al. [50] assume the proximal error bound for $g$ (with $\sigma=0$), equivalent to the global PŁ condition for smooth $g$ [14, Prop. C.1].. The above observations hint at the necessity of a further relaxation to capture richer landscapes.
@@
We relax the global PŁ assumption to the condition, which was recently proposed by [35], to accommodate applications of interest such as those in deep learning. For a fixed $\theta$, a lower-level objective $g(\theta,\cdot)$ admits a provably connected optimal set $\mathcal{S}(\theta)$ and it satisfies the PŁ inequality locally in a neighborhood of $\mathcal{S}(\theta)$. Moreover, the condition ensures that $\mathcal{S}(\theta)$ is a $\mathcal{C}^{2}$ embedding submanifold of the ambient space $\operatorname{\mathbb{R}}^{d}$ without boundary [35]. We shall define condition formally in Section˜2. This setting is sufficiently rich because the lower-level problem admits bounded non-singleton minimizers. Hence, the minima-selection step remains relevant. One such example is when $g$ represents the loss function of some deep learning tasks in the over-parameterized regime, where the local minima are observed to be connected [27, 31, 69, 48, 58]. Another advantage of this setting is that, as we prove in Section˜3.1, the hyper-objective under the condition is Lipschitz continuous. However, under the same condition, the hyper-objective function is not necessarily differentiable, and conditions on functions $f$ and $g$ that would ensure the differentiability of the hyper-objective are often restrictive and unrealistic in practice (see the discussion in Section˜A.2).
@@
Solving BLO problems with a non-singleton lower-level solution set $\mathcal{S}(\theta)$ involves two main challenges: 1) As previously emphasized, for a fixed upper-level variable $\theta$, the minima-selection problem requires maximization over a nonconvex set. Consequently, exact evaluations of the hyper-objective $F_{\max}$ are generally inaccessible. 2) Even when the minima-selection subproblem can be solved exactly, the resulting hyper-objective $F_{\max}$ may still be non-differentiable with respect to the upper-level variable $\theta$. Addressing this issue requires establishing local optimality guarantees for a nonsmooth nonconvex objective, a widely recognized challenge in the literature [46, 44, 78, 47, 93].
To address these issues, we introduce the superquantile–Gibbs relaxation to approximate the hyper-objective, providing a practical surrogate for $F_{\max}$. Using evaluations of this surrogate, we propose a zeroth-order method that seeks a generalized Goldstein stationary point of $F_{\max}$ (see Section˜5.2; cf. [64]).
@@ 1.1 Finding a Goldstein Stationary Point of the Hyper-objective
As we shall see in ˜4.5, under the condition the hyper-objective $F_{\max}$ is Lipschitz continuous. A natural choice for approximate stationarity in this setting is Clarke’s notion of stationarity [18]. However, Zhang et al. [93] showed that, for nonsmooth Lipschitz problems, obtaining an $\epsilon$-Clark stationary point in polynomial time is generally impossible, motivating more nuanced concepts. We therefore adopt the notion of $(\epsilon,\rho)$-Goldstein stationarity [34].
For the domain-constrained objective in (1), akin to [64], we measure stationarity by a generalized $(\epsilon,\rho)$-Goldstein criterion which is defined for the gradient mapping [^3] The gradient mapping for $F:\Theta\to\mathbb{R}$ at a point $\theta$ is $\mathcal{G}_{\Theta}(\theta,\nabla F;\eta):=\eta^{-1}(\theta-\text{Proj}_{\Theta}(\theta-\eta\nabla F(\theta)))$., denoted by $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point (see Section˜5.2). As $F_{\max}$ may not be differentiable, we propose a zeroth-order (ZO) method to find such a stationary point of $F_{\max}$.
@@ 1.2 Superquantile-Gibbs Relaxation (SQ-G)
In this part, we introduce the superquantile of the random variable $f(\theta,X)$ where $X$ is sampled according to the the Gibbs measure $\sim\exp(-g(\theta,x)/\lambda)$. Our idea is that as $\lambda\rightarrow 0$, the Gibbs measure concentrates on the feasible domain $\mathcal{S}(\theta)$, and the superquantile at a high confidence level would approach the maximum value of $f(\theta,\cdot)$ on $\mathcal{S}(\theta)$. Formally, we define the Superquantile-Gibbs Relaxation (SQ-G)-loss using the variational formulation of the superquantile by [74]:
@@ 1.3 Our Contributions
    To allow for lower-level optimal sets with rich structures, we relax the global PŁ condition, commonly imposed on the lower-level objective, to the weaker condition from [35]. We show that the condition is sufficient to ensure Lipschitz continuity of the hyper-objective function $F_{\max}$.
@@
    We use the Superquantile–Gibbs relaxation (SQ-G) to approximate $F_{\max}$. This converts the crucial minima-selection step in BLO, originally a nonconvex optimization over a nonconvex set, into sampling from a Gibbs distribution, which we implement efficiently via Langevin dynamics.
@@
    Assume that $g$ satisfies and the regularity conditions for $f$ and $g$ (to be stated in Section˜2). As shown in Section˜5, by using a Projected Stochastic Zeroth-Order method with Minima-Selection (see Algorithm˜2) to optimize $F_{\max}$, we obtain an $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point of $F_{\max}$ on average using $\tilde{\mathcal{O}}\big(m^{2\texttt{k}+5}\epsilon^{-2}(\epsilon\rho)^{-2\texttt{k}-4}\big)$ samples from the distribution $\mu^{\lambda}_{\theta}$ for $\texttt{k}\geq 1$. When the samples of $\mu^{\lambda}_{\theta}$ are approximated by the Langevin Monte Carlo method, we show in Section˜5.3 that the same goal can be achieved with $\tilde{\mathcal{O}}\big(m^{8\texttt{k}+11}\epsilon^{-2}(\epsilon\rho)^{-8\texttt{k}-10}\big)$ queries to a gradient oracle for $g$ and $\tilde{\mathcal{O}}\big(m^{2\texttt{k}+5}\epsilon^{-2}(\epsilon\rho)^{-2\texttt{k}-4}\big)$ queries to a value oracle for $f$.
@@
    We evaluate our approach on both synthetic and real-world benchmarks and demonstrate the effect of minima-selection. As shown in Section˜6, our proposed approach outperforms hyper-gradient and penalty-based baselines.
@@
(Exponential dependence on k) The $\epsilon^{-\texttt{k}}$ factor in our results reflects the intrinsic difficulty of the problem. Since evaluating $F_{\max}(\theta)=\max_{y\in S(\theta)}f(\theta,y)$ is equivalent to solving a nonconvex optimization problem over a nonconvex set, a single $\epsilon$-accurate evaluation of $F_{\max}(\theta)$ yields an $\Omega(\epsilon^{-\texttt{k}})$ complexity lower bound. More specifically, consider the instance where $S(\theta)=\mathbb{S}^{\texttt{k}}$ is a hypersphere. Reparameterizing $y$ in polar coordinates yields $F_{\max}(\theta)=\max_{x\in\mathbb{B}_{\texttt{k}}}f(\theta,y(x))$. Here $\mathbb{B}_{\texttt{k}}=[0,\pi]^{\texttt{k}}$ and the point $y(x)\in\mathbb{S}^{\texttt{k}}$ is given by $y_{i}(x)=\sin(x_{1})\cdots\sin(x_{i-1})\cos(2x_{i})$ for $i\in[\texttt{k}]$ and $y_{\texttt{k}+1}(x)=\sin(x_{1})\cdots\sin(x_{\texttt{k}-1})\sin(2x_{\texttt{k}})$. It follows from [67, Section 1.6.2] that the worst-case complexity of finding an $\epsilon$-accurate solution to is $\Omega(\epsilon^{-\texttt{k}})$ for $f(\theta,\cdot)\in\mathcal{C}^{1}$.
@@ 1.4 Related Work
Bilevel optimization has a long history in mathematical programming; see, e.g., the early bibliographic review by [81], the monograph [22], the survey [19], the classical text on mathematical programs with equilibrium constraints (MPECs) by [66], and the recent collection [25]. Classical approaches often reduced bilevel problems to single-level programs, either via value-function reformulations or by replacing the lower-level problem with its Karush–Kuhn–Tucker (KKT) optimality system. A related value-function-based reduction for bilevel programs with nonconvex lower-level objectives yields single-level formulations with nonsmooth inequality constraints, for which smoothing and augmented-Lagrangian type algorithms were developed in, e.g., [56, 84]. The latter route yields MPEC formulations in which complementarity conditions between multipliers and constraint violations appear as explicit equilibrium constraints; see, for example, [89, 88, 26, 91]. In practice, these complementarity constraints are frequently handled through mixed-integer or so-called big-$M$ linearizations; however, their performance can be sensitive to the choice of constants, with overly small values excluding feasible bilevel solutions and overly large values leading to weak relaxations and numerical instability [45]. Complementing these algorithmic considerations, recent work has also highlighted inherent geometric and computational hardness in bilevel programming, even under smoothness, which motivates the search for additional structural assumptions and tailored relaxations [12].
@@
Modern large-scale machine-learning applications led to gradient-based methods for bilevel empirical risk minimization and hyperparameter optimization. When the lower-level objective is strongly convex in the inner variable and $f,g$ are smooth, the hyper-objective is differentiable and its gradient can be obtained by implicit differentiation. Representative works in this setting include approximate hyper-gradient methods such as HOAG [70], stochastic bilevel schemes [32, 85], and complexity-refined algorithms for empirical risk minimization [21, 15], as well as more abstract treatments of optimization layers via automatic differentiation [1], amortized implicit differentiation [3], and modular implicit differentiation frameworks [10, 37]. In these approaches, computing exact hyper-gradients typically required Hessian or Hessian-inverse information for the lower-level problem, which could be prohibitive in large neural networks. This motivated a substantial line of Hessian-free or fully first-order methods that approximated the hyper-gradient using only first-order information, including BOME [59], Hessian/Jacobian-free stochastic methods with improved sample complexity [86], fully first-order stochastic approximation methods such as F2BO [49], and bilevel stochastic-gradient methods that also accommodate nonlinear lower-level constraints and inexact hyper-gradient computations [33]. A different line of work revisits bilevel learning from a functional viewpoint: Petrulionytė et al. [72] treat the lower-level variable as a prediction function (rather than a parameter vector) and optimize it over a convex function space. Assuming strong convexity in this functional variable yields a unique lower-level solution and enables implicit differentiation via a functional implicit-function theorem, shifting the strong-convexity requirement from parameters to function space. These developments made bilevel optimization scalable in settings with a unique lower-level solution, so the minima-selection problem that we study did not arise.
@@
A complementary line of work relaxed lower-level strong convexity to a global Polyak–Łojasiewicz (PŁ) condition and developed bilevel gradient methods under this assumption [14, 50, 83, 40, 77, 59]. To obtain a smooth and well-behaved hyper-objective, these methods typically combined the global PŁ condition with additional regularity assumptions on the lower-level problem (as discussed earlier), so that minima-selection was effectively avoided. In contrast, our analysis is based on a local PŁ condition and explicitly exploits the geometry of nontrivial minimizer manifolds. In a complementary direction, Chen et al. [13] introduce a set-smoothness property of the lower-level solution mapping and show that it yields weak convexity/concavity of the induced optimistic/pessimistic hyper-objectives, enabling the computation of approximate Clarke hyper-stationary points. Their analysis relies on assumptions such as a uniform error bound together with smoothness, which imply a global PŁ condition for the lower-level problem. When the lower-level objective is $\mathcal{C}^{2}$ in the lower-level variable, this global PŁ property forces $\operatorname*{arg\,min}_{x}g(\theta,x)$ to be either a singleton or an unbounded set [20]. As a consequence, these conditions rule out the bounded non-singleton minimizer manifolds that motivate the minima-selection regime studied in our work.
Alternative formulations of BLO.